### v1

# Dynamics of the unitary Bose gas near a narrow Feshbach resonance: universal coherent atom-molecule oscillations

Abstract
Quench experiments on a unitary Bose gas around a broad Feshbach resonance have led to the discovery of universal dynamics. This universality is manifested in the measured atomic momentum distributions where, asymptotically, a quasi-equilibrated metastable state is found in which both the momentum distribution and the time scales are determined by the particle density. In this paper we present counterpart studies but for the case of a very narrow Feshbach resonance of 133Cs atoms having a width of 8.3 mG. In dramatic contrast to the behavior reported earlier, a rapid quench of an atomic condensate to unitarity is observed to ultimately lead to coherent oscillations involving dynamically produced condensed and non-condensed molecules and atoms. The same characteristic frequency, determined by the Feshbach coupling, is observed in all types of particles. To understand these quench dynamics and how these different particle species are created, we develop a beyond Hartree-Fock-Bogoliubov dynamical framework including a new type of cross correlation between atoms and molecules. This leads to a quantitative consistency with the measured frequency. Our results, which can be applied to the general class of bosonic superfluids associated with narrow Feshbach resonances, establish a new paradigm for universal dynamics dominated by quantum many-body interactions.
Introduction. Understanding the inherently unstable unitary Bose gas has remained a challenge  Papp et al. (2008); Navon et al. (2011); Pollack et al. (2009); Piatecki and Krauth (2014); Wild et al. (2012); Braaten et al. (2011). Some progress, however, has been made principally because rapid field sweeps across a Feshbach resonance to unitarity show, fortuitously, that the gas lives long enough to reveal features of quasi-steady state behavior before inevitable losses set in. This is seen through the momentum distribution $n(k)$ of non-condensed particles  Makotyn et al. (2014); Eigen et al. (2018). Importantly, these sweeps establish that the signal at high momentum $k$ grows and eventually saturates, as a function of time. These intriguing saturation phenomena have been referred to as a form of prethermalization Yin and Radzihovsky (2016); Rançon and Levin (2014). Interestingly, for the systems studied thus far, the time scale for such quasi-steady state dynamics is claimed to be universal. As the scattering length $a_{s}\rightarrow\infty$, it is presumed that time dependent phenomena should be determined by a single energy scale $E_{F}\propto(6\pi^{2}n)^{2/3}$ Cowell et al. (2002); Song and Zhou (2009); Lee and Lee (2010); Diederix et al. (2011); Li and Ho (2012); Jiang et al. (2014), where $n$ is the density of bosons.
Here we emphasize that this form of universality does not apply to the narrow resonance case. Even at unitarity, the relevant energy scale corresponds to the many-body Feshbach coupling $\alpha\sqrt{n}$, which is to be contrasted with the Fermi energy $E_{\mathrm{F}}$. This scale may be ignored in previous work Makotyn et al. (2014); Eigen et al. (2018) since there $\alpha\sqrt{n}\gg E_{F}$. However, the dynamics of the unitary gas in this narrow resonance regime which has not received the same degree of attention, remains to be experimentally characterized and theoretically understood.
Figure 1: Effective temperature of atoms (black solid line) and molecules (blue) based on an error-bar weighted fit to data from Ref.Zhang et al. (2023) emphasizing the creation of out-of-condensate particles deriving from an atomic BEC after a quench to unitarity, see Supplement sup . The small cartoons indicate the different species of particles generated at different stages of the dynamics, including atomic and molecular BECs, finite-momentum atomic and molecular pairs.
In this paper, using both theory and experiment, we investigate the case of this narrow Feshbach resonance. In contrast to previous Bose gas literature which focused on the atoms, one has access to a sizeable population of closed channel molecules which govern much of the physics at unitarity. Our work is motivated by an earlier observation of the effective temperature, $T_{\rm{eff}}$ of out-of-condensate atoms and of closed channel molecules Zhang et al. (2023) which is plotted in Fig. 1. These are created during a quench to unitarity of an atomic Bose-Einstein condensate (BEC) of 133Cs atoms associated with a narrow resonance (of width 8.3 mG). Although previously oscillations in $T_{\rm{eff}}$ were not reported, they are revealed here based on an error-bar weighted fit to the temperature data.
It is our goal to experimentally investigate and theoretically understand the initial creation and subsequent oscillatory dynamics of these excited atomic and molecular states. To this end, we employ a time-of-flight (TOF) procedure. By analyzing a data set from previous experiments Zhang et al. (2023); Wang et al. (2024), we are able to characterize multiple new observables. These include the atomic and molecular particle number distribution at different momenta and their associated kinetic energies sup .
Theoretically to understand the oscillatory behavior observed in these quantities, we go beyond the Hartree-Fock-Bogolibov (HFB)  Kokkelmans and Holland (2002); Snyder et al. (2012) approach and establish a new dynamical framework that includes atom-molecule correlations. These correlations, which have no counterpart in the broad resonance case, turn out to be central to the dynamics of the out-of-condensate particles. In the process we identify three distinct evolutionary stages in the quench dynamics; this is initiated by the generation of a molecular BEC and subsequently followed by the creation of finite-momentum atoms and molecules which then undergo coherent oscillationsZhang et al. (2023); Tian et al. (2024). Satisfactory quantitative agreement is obtained between theory and experiment both for the rather large oscillatory frequency which turns out to be around $2$ kHz and for the net increase of kinetic energy (associated with the decay of the initial atomic condensate) which is transferred to newly excited atoms and molecules.
Dynamical Framework. The effective Hamiltonian for a magnetic Feshbach resonance is associated with coupling between atomic and molecular channels (respectively called “open” and “closed”). It is given by
$\hat{H}=\sum_{\sigma}\int d^{3}x\hat{\psi}^{\dagger}_{\sigma}(x)\left(-\frac{\hbar^{2}\nabla^{2}}{2m_{\sigma}}+\nu_{\sigma}\right)\hat{\psi}_{\sigma}(x)$ (1)
$+\int d^{3}x\sum_{\sigma}\frac{g_{\sigma}}{2}\hat{\psi}^{\dagger}_{\sigma}\hat{\psi}^{\dagger}_{\sigma}\hat{\psi}_{\sigma}\hat{\psi}_{\sigma}-\left(\alpha\hat{\psi}^{\dagger}_{1}\hat{\psi}^{\dagger}_{1}\hat{\psi}_{2}+h.c.\right).$
Here $\sigma=1,2$ represents the atoms and molecules, respectively, and $\nu_{\sigma}=\delta_{\sigma,2}\nu$ where $\nu$ is the detuning; $g_{\sigma}$ is proportional to the background $s$-wave interaction, and $\alpha$ corresponds to the important Feshbach coupling strength between the two channels. Model parameters are related to the physical quantities by $g_{\sigma}=4\pi\hbar^{2}a_{\sigma}/(\Gamma_{\sigma}m_{\sigma})$, $\alpha=\Gamma_{1}^{-1}\sqrt{2\pi\hbar^{2}a_{1}\delta\mu\Delta B/m_{1}}$ and $\Gamma_{\sigma}$ is $1-2a_{\sigma}\Lambda/\pi$ where $\Lambda$ is a regularization cut-off. Here $a_{1}$ and $a_{2}$ are the background scattering length of atom and molecules, $\delta\mu$ is the relative magnetic moment between open and closed channels and $\Delta B$ is the resonance width. The detuning $\nu$ is related to the physical detuning $\nu_{r}=\delta\mu\delta B$ by $\nu=\nu_{r}+\Lambda m_{1}\Gamma_{1}\alpha^{2}/(\pi^{2}\hbar^{2})$ where $\delta B$ is the distance of the magnetic field from the resonance. We use the experimental values Zhang et al. (2023) for the parameters $a_{1},a_{2},\delta\mu,\Delta B$ and the initial atomic density $n$ in the theoretical calculations sup .
To theoretically study non-equilibrium dynamics, we focus on one and two-point equal-time correlations as dynamical variables. This approach is related to others in the literature Colussi et al. (2020); Van Regemortel et al. (2018) but now generalized to include the complexity of the two channel system. The one-point correlation $\xi_{\sigma}(t)=V^{-1/2}\langle\hat{\psi}_{\sigma}(k=0,t)\rangle$ represents the atomic or molecular condensate wavefunction, and $c_{\sigma}\equiv|\xi_{\sigma}(t)|^{2}$ represents the condensate density. Operators representing the excitations are then given by $\hat{\psi}_{\sigma}^{\prime}(k)=\hat{\psi}_{\sigma}(k)-\delta_{k,0}\sqrt{V}\xi_{\sigma}(t)$.
It is convenient to introduce a 4-vector field operator, $\hat{\Psi}(k)~{}=[\hat{\psi}^{\prime}_{1}(k),\hat{\psi}^{\prime\dagger}_{1}(-k),\hat{\psi}^{\prime}_{2}(k),\hat{\psi}^{\prime\dagger}_{2}(-k)]$. The quantum fluctuations that involve the finite momentum particles correspond to two-point correlation functions, $G_{\alpha\beta}(k)=\langle\hat{\Psi}_{\alpha}(-k)\hat{\Psi}_{\beta}(k)\rangle$, from which we deduce physical quantities. The variables $n_{1}(k)\equiv G_{21}(k)$ and $n_{2}(k)\equiv G_{43}(k)$ represent the particle number associated with finite-$k$ atoms and molecules. Important are the interspecies correlations $G_{13}(k)$ and $G_{23}(k)$ which are central to the dynamics: $G_{13}(k)$ is associated with the molecular pairs and $G_{23}(k)$ controls the particle flow which results in the creation of finite-$k$ molecules.
The dynamical equations associated with these correlation functions are readily derived as
$i\partial_{t}\xi_{1}=2g_{1}n_{1}\xi_{1}+\left(g_{1}\xi_{1}^{2}+g_{1}x_{1}-2\alpha\xi_{2}\right)\xi^{*}_{1}-2\alpha f_{12},$ (2)
$i\partial_{t}\xi_{2}=\left(\nu+2g_{2}n_{2}\right)\xi_{2}+(g_{2}\xi_{2}^{2}+g_{2}x_{2})\xi^{*}_{2}-\alpha(x_{1}+\xi_{1}^{2}).$
Here $n_{\sigma}=\sum_{k\neq 0}n_{\sigma}(k)/V$ is the density of out-of-condensate atoms and molecules and $V$ is the volume. We define $x_{1}=\sum_{k\neq 0}G_{11}(k)/V$ and $x_{2}=\sum_{k\neq 0}G_{33}(k)/V$ which represent atom and molecule pairs, respectively, as discussed in the supplement sup . Additionally $f_{12}=\sum_{k}G_{23}(k)/V$ characterizes the atom-molecule correlation function which plays an important role. The equations of motion for the 2-point correlations can be compactly written as
$i\partial_{t}G_{mn}(k)=\sum_{\beta}L_{m\beta}(k)G_{\beta n}(k)+L_{n\beta}(k)G_{m\beta}(k)$ (3)
where $L(k)$ represents the coupling between finite-momentum atoms and molecules sup . This dynamical scheme should be contrasted with the standard Hartree-Fock-Bogoliubov approach Kokkelmans and Holland (2002); Snyder et al. (2012) which omit these new atom-molecule correlations including $f_{12}$.
Figure 2: Population transfers and growth in the present theory. When an atomic BEC is quenched to unitarity, a molecular BEC is the first to be created. After $t=0.2$ ms, both atomic and molecular BECs decay, generating a large fraction of non-condensed atoms and molecules. After 1 ms, all particles exhibit oscillations with identical frequency. Only the non-condensed molecules are out of phase as they control the flow of particles.
Figure 3: Kinetic energy variation $E_{K}/N_{0}$ for atoms and molecules which is associated with the population transfers and growth in Figure 2. Here $E^{\sigma}_{K}/N_{0}\equiv\int d^{3}k/(2\pi)^{3}n_{\sigma}(k)(k_{x}^{2}+k_{y}^{2})/(2m_{\sigma}n)$. Theory (top panel) and experiment (bottom panel). In experiment (but not theory) there is loss of particles which is associated with a gain in kinetic energy, but the two effects tend to at least partially compensate in the figure. Experiment and theory show similar oscillatory behavior and both show a net increase of kinetic energy which is of the order of $5$ nK times $N_{0}$, where $N_{0}$ represents the total number of particles at the start of the sweep.
Three Evolutionary Stages Using these dynamical equations we identify different stages of particle creation throughout the quench dynamics. The quench of an atomic BEC to unitarity is implemented by an abrupt change in the detuning $\nu_{r}\sim 0$ assuming initially, at $t=0$, all particles reside in an atomic BEC. Fig. 2 indicates the three distinct dynamical stages where different species of particles are generated. In stage I, Feshbach couplings convert the initial atomic condensate into a coherent molecular BEC along with a small number of atomic and molecular pairs having momentum, ±$\pm k$. 111 The growth of this molecular BEC is associated with instanton dynamics: $\xi_{2}(t)/\sqrt{n}=i\tanh(t/t_{\alpha})$, where $t_{\alpha}=\hbar/\alpha\sqrt{2n}$ sup . In stage II, shortly after its formation, the molecular BEC begins to partially decay and more atomic pairs appear. These latter in turn combine with an atom from the condensate to form an atom-molecule complex. The combination of an atom-molecule complex and an atom from the condensate is further converted into a pair of molecules. In stage III, with these finite-momentum atomic and molecular populations now fully formed, the system enters a stage of steady-state coherent oscillations.
What is particularly interesting about these oscillations is that, at unitarity, there is an in-phase relationship between all quantities except for the non-condensed molecules. This phenomenon derives from a competition involving the three particle currents driven by the Feshbach coupling sup . These currents represent the flow of particles between atoms and molecules and condensates. Our central finding is that at resonance, once non-condensed molecules are created, they tend to control (through the equations of motion) the general dynamics of the particles. As a consequence, the out-of-condensate molecules are, interestingly, out-of-phase with the other three species sup .
Figure 4: Theory-experiment comparison of the dynamics of finite-momentum molecules where (a) represents experimentally measured momentum distribution $f_{m}$ at 3 different times showing non-monotonicity– the ordering of the curves from top to bottom corresponds to early, then late and then middle time. Here $f_{m}$ is the angular average momentum distribution of molecules such that $\int d^{2}kf_{m}=N_{m}$. (b) plots their number $vs$ time, $t$, when summed over a momentum shell. The summation region is indicated by $p_{1}/p_{2}/p_{3}$. Here $p_{i}=\int_{k_{i}\leq k<k_{i+1}}d^{2}kf_{m}(k)$ where $k_{i}=0.12k_{F}\times i$. Panels (c) and (d) indicate the counterpart theory plots. We observe coherent oscillations in three different shells all with the same frequency of 2kHz in both theory and experiment. Comparing panels (a) and (c) suggests that the occupation of more and higher momentum states seen in experiment may reflect thermal effects: the experiments are performed at finite $T$, whereas theory is at $T=0$.
Most of the general observations we outline below (unless indicated otherwise) pertain to both theory and experiment 222 Because the experimentally observed particle loss is not significant within the 3 $\mathrm{ms}$ time frame Zhang et al. (2023), it makes sense to consider these comparisons. and these can be seen in Fig. 3. As is suggested by Figure 1, we find the newly generated out-of-condensate particles carry kinetic energy that also undergoes oscillatory dynamics. As the atomic condensate decays, a portion of the interaction energy is transferred to non-condensed particles as kinetic energy. This leads to an increase of the kinetic energy in both atoms and molecules right after the quench at $t<0.5~{}\mathrm{ms}$. We can quantify this increase, $E^{\sigma}_{K}/N_{0}$, which is of the order of 10 $\mathrm{nK}$ for the atoms and 5 $\mathrm{nK}$ for the molecules, as found in both theory and experiment.
After this initial rise, the kinetic energy subsequently oscillates and these oscillations persist for a period of around 2 $\mathrm{ms}$ notably, without any sign of thermalization. This oscillatory frequency is around 2 kHz, which, importantly, coincides with that of the population oscillations (see Fig. 2Zhang et al. (2023).) Indeed, this single oscillation pattern observed in all collective observables reflects the fact that the same underlying microscopic dynamical processes are present.
Momentum-resolved Distributions and Asymptotic Coherent Oscillations. Of interest next is to clarify the microscopic mechanism responsible for the generation of finite-momentum particles, after the initial formation of the molecular BEC. This is associated with the Feshbach resonant interaction and is contained in the properties of the matrix $L(k)$ in Eq. 3. The eigenvalues of $L$ determine the speed of growth of these out-of-condensate particles while the eigenvectors determine the fraction of atoms and molecules involved in the generation process. We consider the eigenvalues of $L(k)$,
$\lambda_{\pm}\simeq\sqrt{F_{1}(k)\pm\sqrt{F^{2}_{1}(k)-4F_{2}(k)}}.$ (4)
Here $F_{1}/F_{2}$ are given by $F_{1}(k)=\epsilon^{2}_{1}(k)+\epsilon^{2}_{2}(k)+8\alpha^{2}|\xi_{1}|^{2}-4\alpha^{2}|\xi_{2}|^{2}$ and $F_{2}(k)=(\epsilon_{1}(k)\epsilon_{2}(k)-4\alpha^{2}|\xi_{1}|^{2})^{2}-4\alpha^{2}|\xi_{2}|^{2}\epsilon^{2}_{2}(k)$. This depends on dispersions given by $\epsilon_{\sigma}(k)=k^{2}/2m_{\sigma}+\nu_{r}\delta_{2,\sigma}$. One may observe that there exist two circumstances which lead to an imaginary $\lambda$ and a consequent instability of condensates: either $F_{1}(k)<0$ which is associated with the processes in stage I, or $F_{2}(k)<0$ which is mainly associated with those processes in stage II. These unstable outcomes are unavoidable when particles mostly reside in condensates. Indeed, once a molecular condensate is present ( $\xi_{2}\neq 0$), one can always find a value of $k$ such that $\epsilon_{1}\epsilon_{2}-4\alpha^{2}|\xi_{1}|^{2}\simeq 0$ leading to an imaginary value for $\lambda_{-}$. Additionally, for $|\xi_{2}|\gg|\xi_{1}|$, the instability occurs at $\epsilon^{2}_{1}(k)+\epsilon^{2}_{2}(k)<4\alpha^{2}|\xi_{2}|^{2}$. In this way, finite-momentum particles are exponentially generated.
All of this can be contrasted with the larger $k$ regime where the eigenvalues are real and there is consequently no significant occupation of higher momentum states, say $k$ of the order of $k_{\mathrm{F}}$. Here the kinetic oscillations rather than the exponential growth dominates. Indeed, the momentum distributions we observe are consistent with the above analysis, where particles are seen to be concentrated in states with low momentum up to $\approx k_{F}/2$ where $k_{F}\equiv(6\pi^{2}n)^{1/3}$. We stress that, in contrast to the saturation effects seen in the broad Feshbach resonanceMakotyn et al. (2014); Eigen et al. (2018), the momentum distributions in the narrow resonance case oscillate during the later time dynamics, as seen in Fig. 4a,c.
To confirm that a single oscillation frequency is present in the experiments, as suggested by theory, we determine the molecular population in a shell including different groupings of momenta $p_{i}$, as in Fig. 4. Three such subsets were considered and the behavior in all is found to be rather similar with all showing the characteristic frequency of $2$ kHz. This shared common frequency between finite-momentum atoms and molecules and their condensates should be understood to derive from Feshbach coupling sup . We note parenthetically that the amplitude of the oscillations is slightly greater in theory than experiment, which we attribute to finite temperature effects in our experiment that suppress the coherence of atoms and molecules.
[…]
Conclusion. We have emphasized throughout that the coherent oscillations we observe for the narrow resonance case are universal as they are only determined by a single energy scale $\alpha\sqrt{n}$. This leads to a different paradigm for the universal dynamics of the unitary Bose gas. This is associated with a many body (Feshbach) coupling and is to be contrasted with the universality previously discussed in the literature  Makotyn et al. (2014); Eigen et al. (2018) and associated with the Fermi energy. As in this early literature, arriving at this universality requires us to focus on characterizing the dynamical behavior of non-condensed particles which appear when an atomic condensate is swept to unitarity. Here in addition to atoms there are out-of-condensate molecules, and once formed, all populations reach a steady oscillatory state; all finite-momentum particles are seen to have the same intrinsic oscillation frequency which is the same as that of the condensate Wang et al. (2024).
In this context, we have presented a reasonably successful comparison between theory and experiment. Because of these quantitative correspondences we are able to glean support for a rather new and different dynamical theory. This is to be contrasted with the more conventional Hartree-Fock-Bogoliubov approach  Kokkelmans and Holland (2002); Snyder et al. (2012). Importantly, this dynamical machinery and its rich phenomenology should be relevant to the general class of bosonic superfluids associated with narrow Feshbach resonances.
Acknowledgement This work is supported by the National Science Foundation under Grant No. PHY1511696 and PHY-2103542, and by the Air Force Office of Scientific Research under award number FA9550-21-1-0447. We thank Cheng Chin for helpful discussions. Z.Z. acknowledges the Bloch Postdoctoral Fellowship. S.N. acknowledges support from the Takenaka Scholarship Foundation. Z.W. is supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.